# Mirror Descent Policy Optimisation for Robust Constrained Markov Decision Processes

### v2

@@ C.2 Regret bound for approximate TMA
$\langle p^{*}(\cdot|s,a)-p^{t}(\cdot|s,a),G_{\pi,p^{t}}(s,a,\cdot)\rangle+\langle p^{t}(\cdot|s,a)-p^{t+1}(\cdot|s,a),G_{\pi,p^{t}}(s,a,\cdot)\rangle\leq\frac{\alpha_{p}(t)}{\eta_{p}(t)}\left(B(p^{*},p^{t})-B(p^{*},p^{t+1})\right)$
@@
$\underbrace{\frac{1}{1-\gamma}\sum_{s\in\mathcal{S}}d_{\rho}^{\pi,p^{*}}(s)\sum_{a\in\mathcal{A}}\pi(a|s)\langle p^{*}(\cdot|s,a)-p^{t}(\cdot|s,a),\hat{G}_{\pi,p^{t}}(s,a,\cdot)\rangle}_{A}+$
$\underbrace{\frac{1}{1-\gamma}\sum_{s\in\mathcal{S}}d_{\rho}^{\pi,p^{*}}(s)\sum_{a\in\mathcal{A}}\pi(a|s)\langle p^{t}(\cdot|s,a)-p^{t+1}(\cdot|s,a),\hat{G}_{\pi,p^{t}}(s,a,\cdot)\rangle}_{B}$
@@
$\frac{1}{1-\gamma}\sum_{s\in\mathcal{S}}d_{\rho}^{\pi,p^{*}}(s)\sum_{a\in\mathcal{A}}\pi(a|s)\left(\langle p^{*}(\cdot|s,a)-p^{t}(\cdot|s,a),G_{\pi,p^{t}}(s,a,\cdot)\rangle+\langle p^{*}(\cdot|s,a)-p^{t}(\cdot|s,a),\hat{G}_{\pi,p^{t}}(s,a,\cdot)-G_{\pi,p^{t}}(s,a,\cdot)\rangle\right)$ (decomposing)
$\geq V_{\pi,p^{*}}(\rho)-V_{\pi,p^{t}}(\rho)+\frac{1}{1-\gamma}\sum_{s\in\mathcal{S}}d_{\rho}^{\pi,p^{*}}(s)\sum_{a\in\mathcal{A}}\pi(a|s)\langle p^{*}(\cdot|s,a)-p^{t}(\cdot|s,a),\hat{G}_{\pi,p^{t}}(s,a,\cdot)-G_{\pi,p^{t}}(s,a,\cdot)\rangle$ (Lemma 22)
$\geq V_{\pi,p^{*}}(\rho)-V_{\pi,p^{t}}(\rho)-\frac{2C\epsilon^{\prime}}{1-\gamma}\,$,
where the last step bounds the absolute value as in the proof of Lemma 5.
For part B), note that
$\frac{1}{1-\gamma}\sum_{s\in\mathcal{S}}d_{\rho}^{\pi,p^{*}}(s)\sum_{a\in\mathcal{A}}\pi(a|s)\langle p^{t}(\cdot|s,a)-p^{t+1}(\cdot|s,a),\hat{G}_{\pi,p^{t}}(s,a,\cdot)\rangle$
@@
$\frac{1}{1-\gamma}\sum_{s\in\mathcal{S}}\frac{d_{\rho}^{\pi,p^{*}}(s)}{d_{\rho}^{\pi,p^{t+1}}(s)}d_{\rho}^{\pi,p^{t+1}}(s)\sum_{a\in\mathcal{A}}\pi(a|s)\big(\langle p^{t}(\cdot|s,a)-p^{t+1}(\cdot|s,a),G_{\pi,p^{t}}(s,a,\cdot)\rangle+$
@@
$\hskip 113.81102pt\frac{1}{1-\gamma}\sum_{s\in\mathcal{S}}d_{\rho}^{\pi,p^{t+1}}(s)\sum_{a\in\mathcal{A}}\pi(a|s)\langle p^{t}(\cdot|s,a)-p^{t+1}(\cdot|s,a),\hat{G}_{\pi,p^{t}}(s,a,\cdot)-G_{\pi,p^{t}}(s,a,\cdot)\rangle\Big)$ (ascent property (Eq. 28))
@@
$\hskip 113.81102pt\frac{1}{1-\gamma}\sum_{s\in\mathcal{S}}d_{\rho}^{\pi,p^{t+1}}(s)\sum_{a\in\mathcal{A}}\pi(a|s)\langle p^{t}(\cdot|s,a)-p^{t+1}(\cdot|s,a),\hat{G}_{\pi,p^{t}}(s,a,\cdot)-G_{\pi,p^{t}}(s,a,\cdot)\rangle\Big)$ (definition of $M=\sup_{p\in\mathcal{P}}\left\lVert\frac{d_{\rho}^{\pi,p^{*}}}{d_{\rho}^{\pi,p}}\right\rVert_{\infty}$ and again Eq. 28 )
$=M\Big(\left(V_{\pi,p^{t}}(\rho)-V_{\pi,p^{t+1}}(\rho)\right)+$
$\hskip 113.81102pt\frac{1}{1-\gamma}\sum_{s\in\mathcal{S}}d_{\rho}^{\pi,p^{t+1}}(s)\sum_{a\in\mathcal{A}}\pi(a|s)\langle p^{t}(\cdot|s,a)-p^{t+1}(\cdot|s,a),\hat{G}_{\pi,p^{t}}(s,a,\cdot)-G_{\pi,p^{t}}(s,a,\cdot)\rangle\Big)$ (first performance difference lemma (Lemma 22))
@@ C.3 Approximate action next-state value
$|\mathbb{E}\left[\gamma^{t}c(s,a,s^{\prime})\right]|\leq\sum_{t=N_{G,t}}^{\infty}C\gamma^{t}=C\frac{\gamma^{N_{G,t}}}{1-\gamma}\,.$
@@ D.1 Analysis of the multipliers
To analyse the regret in terms of the value directly, the lemma below provides an equivalence between the unconstrained value and the Lagrangian value at the optimum.
@@
Lemma 14 (Lower bound on the inner product, Eq. 41 in Liu et al., 2021b ).
@@ D.2 Analysis of the Bregman divergence
Another essential part of the regret analysis is the pushback property, which allows a telescoping sum.
Lemma 15 (Pushback property, Lemma 2 in Liu et al., 2021b ).
@@
Lemma 16 (Pushback property, Lemma 11 in Liu et al., 2021b rephrased).
@@ D.3 Convergence of approximate entropy-regularised NPG
Lemma 18 (Convergence of approximate entropy-regularised NPG, Theorem 2 of Cen et al., (2022)).
@@
Lemma 19 (Number of inner loop iterations, Lemma 7 in Liu et al., 2021b ).
@@
Lemma 20 (Performance difference lemma, Lemma 4 in Cen et al., (2022) and Lemma 6 in Liu et al., 2021b ).
@@ D.4 Performance difference across transition kernels
Lemma 21 (First performance difference lemma across transition kernels, Lemma 5.2 in Wang et al., (2023)).
@@
Lemma 22 (Second performance difference lemma across transition kernels, Lemma 5.3 in Wang et al., (2023)).
@@ D.6.1 Supporting lemmas
The following lemma provides an upper bound on the regret difference induced by the LTMA update.
Lemma 23 (Analysis of the regret difference).
Let $\pi^{*}\in\mathrm{arg}\min_{\Pi}\Phi(\pi)$. Moreover, let $\mathcal{P}$ be an uncertainty set contained in an $\ell_{q}$ $(s,a)$-rectangular uncertainty set around the nominal distribution $\bar{p}=p_{0}$ such that
@@
for some $q\geq 1$, where $\Delta_{p}=\max_{s,a}\psi_{s,a}=\mathcal{O}\left(\frac{S^{1/2+g(q)-1}(1-\gamma)^{2}}{A}\right)$ for $g(q)=\max\{0,\frac{1}{2}-\frac{1}{p}\}$. Then the difference in cumulative regret induced by the transition kernel updates in Algorithm 2 is bounded by a constant, i.e.
$\sum_{k=0}^{K-1}V_{\pi_{k+1},p_{k+1}}(\rho)-V_{\pi^{*},p_{k+1}}(\rho)-\left(V_{\pi_{k+1},p_{k}}(\rho)-V_{\pi^{*},p_{k}}(\rho)\right)=\mathcal{O}(1)\,.$
@@
$\leq\frac{2}{(1-\gamma)^{2}}SAS^{-(1/2+g(p))}\max_{s\in\mathcal{S},a\in\mathcal{A}}\left\lVert p_{K}(\cdot|s,a)-p_{0}(\cdot|s,a)\right\rVert_{p}$ (for any $X\in\mathbb{R}^{d}$, $\left\lVert X\right\rVert_{1}\leq d^{1/2}\left\lVert X\right\rVert_{2}\leq d^{1/2+g(p)}\left\lVert X\right\rVert_{p}$ via Hölder’s inequality)
$\leq\frac{2}{(1-\gamma)^{2}}AS^{-(1/2+g(p)-1)}\Delta_{p}$ (definition of $\Delta_{p}$)
$=\mathcal{O}(1)\,$, (65)
where the last step follows from $\Delta_{p}=\mathcal{O}\left(\frac{S^{1/2+g(q)-1}(1-\gamma)^{2}}{A}\right)$. ∎
@@
Lemma 25 (Lemma 16 of Liu et al., 2021b rephrased).
@@
For any $\pi\in\Pi,p\in\mathcal{P}$, we have $d_{\rho}^{\pi,p}(s)\geq(1-\gamma)\rho(s)$, such that using derivations as in Theorem 4.2 of Wang et al., (2024),
@@ D.6.2 Main theorem
(a) The proof uses similar derivations as in Theorem 3 of Liu et al., 2021b and then accounts for the performance difference due to transition kernels.
@@
Filling in the lower bound on the inner product from Eq. 41 from Liu et al., 2021b (see also Lemma 14),
@@
Applying Lemma 25 to $p^{*}$, it follows that $\left\lVert\lambda^{*}\right\rVert=\mathcal{O}(1)$ such that $\sum_{j=1}^{m}\lambda_{j}^{*}\mathcal{O}(\epsilon)=\mathcal{O}(\epsilon)$. Note then that via Lemma 24, $\frac{\alpha}{1-\gamma}\sum_{k\in[K]}\left(B_{d_{\rho}^{\pi^{*},p_{k-1}}}(\pi^{*},\pi_{k-1})-B_{d_{\rho}^{\pi^{*},p_{k-1}}}(\pi^{*},\pi_{k})\right)\leq\frac{\alpha}{1-\gamma}\left(\log(A)+2\log(\frac{1}{U})\right)$. Moreover, the preconditions for $\alpha$ in the further derivations in Lemma 21 of Liu et al., 2021b are satisfied. The remainder of the proof removes the constant $\frac{M}{1-\gamma}=\mathcal{O}(1)$ and then follows the derivations in Lemma 21 of Liu et al., 2021b . ∎

### v4

@@ C.2 Regret bound for approximate TMA
$\langle p^{*}(\cdot|s,a)-p^{t}(\cdot|s,a),\hat{G}_{\pi,p^{t}}(s,a,\cdot)\rangle+\langle p^{t}(\cdot|s,a)-p^{t+1}(\cdot|s,a),\hat{G}_{\pi,p^{t}}(s,a,\cdot)\rangle\leq\frac{\alpha_{p}(t)}{\eta_{p}(t)}\left(B(p^{*},p^{t})-B(p^{*},p^{t+1})\right)$
@@
$\underbrace{\frac{1}{1-\gamma}\sum_{s\in\mathcal{S}}d_{\rho}^{\pi,p^{*}}(s)\sum_{a\in\mathcal{A}}\pi(a|s)\langle p^{*}(\cdot|s,a)-p^{t}(\cdot|s,a),\hat{G}_{\pi,p^{t}}(s,a,\cdot)\rangle}_{A}$
$+\underbrace{\frac{1}{1-\gamma}\sum_{s\in\mathcal{S}}d_{\rho}^{\pi,p^{*}}(s)\sum_{a\in\mathcal{A}}\pi(a|s)\langle p^{t}(\cdot|s,a)-p^{t+1}(\cdot|s,a),\hat{G}_{\pi,p^{t}}(s,a,\cdot)\rangle}_{B}$
@@
$=\frac{1}{1-\gamma}\sum_{s\in\mathcal{S}}d_{\rho}^{\pi,p^{*}}(s)\sum_{a\in\mathcal{A}}\pi(a|s)\left(\langle p^{*}(\cdot|s,a)-p^{t}(\cdot|s,a),G_{\pi,p^{t}}(s,a,\cdot)\rangle+\langle p^{*}(\cdot|s,a)-p^{t}(\cdot|s,a),\hat{G}_{\pi,p^{t}}(s,a,\cdot)-G_{\pi,p^{t}}(s,a,\cdot)\rangle\right)$ (decomposing)
$=V_{\pi,p^{*}}(\rho)-V_{\pi,p^{t}}(\rho)+\frac{1}{1-\gamma}\sum_{s\in\mathcal{S}}d_{\rho}^{\pi,p^{*}}(s)\sum_{a\in\mathcal{A}}\pi(a|s)\langle p^{*}(\cdot|s,a)-p^{t}(\cdot|s,a),\hat{G}_{\pi,p^{t}}(s,a,\cdot)-G_{\pi,p^{t}}(s,a,\cdot)\rangle$ (Lemma 23)
$\geq V_{\pi,p^{*}}(\rho)-V_{\pi,p^{t}}(\rho)-\frac{2\epsilon^{\prime}}{1-\gamma}\,$,
For part B), note that applying the ascent property (Eq. 31 on $\hat{G}$ results in
$0$ $\geq\frac{1}{1-\gamma}\sum_{s\in\mathcal{S}}d_{\rho}^{\pi,p^{*}}(s)\sum_{a\in\mathcal{A}}\pi(a|s)\langle p^{t}(\cdot|s,a)-p^{t+1}(\cdot|s,a),\hat{G}_{\pi,p^{t}}(s,a,\cdot)\rangle$
@@
$=\frac{1}{1-\gamma}\sum_{s\in\mathcal{S}}\frac{d_{\rho}^{\pi,p^{*}}(s)}{d_{\rho}^{\pi,p^{t+1}}(s)}d_{\rho}^{\pi,p^{t+1}}(s)\sum_{a\in\mathcal{A}}\pi(a|s)\big(\langle p^{t}(\cdot|s,a)-p^{t+1}(\cdot|s,a),G_{\pi,p^{t}}(s,a,\cdot)\rangle+$
@@
$\hskip 113.81102pt\frac{1}{1-\gamma}\sum_{s\in\mathcal{S}}d_{\rho}^{\pi,p^{t+1}}(s)\sum_{a\in\mathcal{A}}\pi(a|s)\langle p^{t}(\cdot|s,a)-p^{t+1}(\cdot|s,a),\hat{G}_{\pi,p^{t}}(s,a,\cdot)-G_{\pi,p^{t}}(s,a,\cdot)\rangle\Big)$ (since the term is negative, upper bound the ratio; Lemma 23)
@@
$\hskip 113.81102pt\frac{1}{1-\gamma}\sum_{s\in\mathcal{S}}d_{\rho}^{\pi,p^{t+1}}(s)\sum_{a\in\mathcal{A}}\pi(a|s)\langle p^{t}(\cdot|s,a)-p^{t+1}(\cdot|s,a),\hat{G}_{\pi,p^{t}}(s,a,\cdot)-G_{\pi,p^{t}}(s,a,\cdot)\rangle\Big)$ (definition of $M=\sup_{p\in\mathcal{P}}\left\lVert\frac{d_{\rho}^{\pi,p^{*}}}{d_{\rho}^{\pi,p}}\right\rVert_{\infty}$)
@@
where the last step again follows from derivations in Lemma 5.
@@ C.3 Approximate action next-state value
$\left|\mathbb{E}\left[\sum_{t=N_{G,t}}^{\infty}\gamma^{t}c(s_{t},a_{t},s_{t+1})\right]\right|\leq\sum_{t=N_{G,t}}^{\infty}C\gamma^{t}=C\frac{\gamma^{N_{G,t}}}{1-\gamma}\,.$
@@ D.1 Analysis of the multipliers
To analyse the average regret in terms of the value directly, the lemma below provides an equivalence between the unconstrained value and the Lagrangian value at the optimum.
@@
Lemma 15 (Lower bound on the inner product, Eq. 41 in Liu et al. (2021b)).
@@ D.2 Analysis of the Bregman divergence
Another essential part of the average regret analysis is the pushback property, which allows a telescoping sum.
Lemma 16 (Pushback property, Lemma 2 in Liu et al. (2021b)).
@@
Lemma 17 (Pushback property, Lemma 11 in Liu et al. (2021b) rephrased).
@@ D.3 Convergence of approximate entropy-regularised NPG
Lemma 19 (Convergence of approximate entropy-regularised NPG, Theorem 2 of Cen et al. (2022)).
@@
Lemma 20 (Number of inner loop iterations, Lemma 7 in Liu et al. (2021b)).
@@
Lemma 21 (Performance difference lemma, Lemma 4 in Cen et al. (2022) and Lemma 6 in Liu et al. (2021b)).
@@ D.4 Performance difference across transition kernels
Lemma 22 (First performance difference lemma across transition kernels, Lemma 5.2 in Wang et al. (2023)).
@@
Lemma 23 (Second performance difference lemma across transition kernels, Lemma 5.3 in Wang et al. (2023)).
@@ D.6.1 Supporting lemmas
The following lemma provides a constant upper bound on the cumulative regret difference induced by the LTMA update under general uncertainty sets (including non-rectangular uncertainty sets), where constant indicates that it is independent of the number of iterations. With additional $\ell_{q}$ norm bound assumptions on the uncertainty set, the constant can be reduced in rectangular and non-rectangular uncertainty sets. The improvement factor on the upper bound is $\frac{2S^{1-c(q)}}{\Delta_{p}}$ for rectangular $\ell_{q}$ sets and $\frac{2(SA^{2})^{1-c(q)}}{\Delta_{p}}$ for non-rectangular $\ell_{q}$ sets, where c(q)=11/qc(q)=1-1/q and $\Delta_{p}$ is the uncertainty budget (upper bound).
Lemma 24 (Analysis of the cumulative regret difference).
Let $\pi^{*}\in\mathrm{arg}\min_{\pi\in\Pi}\Phi(\pi)$, and define $D\geq\sum_{k=0}^{K-1}V_{\pi_{k+1},p_{k+1}}(\rho)-V_{\pi^{*},p_{k+1}}(\rho)-\left(V_{\pi_{k+1},p_{k}}(\rho)-V_{\pi^{*},p_{k}}(\rho)\right)$ as the difference in cumulative regret induced by the transition kernel updates in Algorithm 3.
a) For any uncertainty set $\mathcal{P}$, $D\leq\frac{4}{(1-\gamma)^{2}}AS^{2}$.
b) Let $\mathcal{P}$ be an uncertainty set contained in an $\ell_{q}$ $(s,a)$-rectangular uncertainty set around the nominal $\bar{p}=p_{0}$ such that
@@
for some $q\geq 1$, where $0<\Delta_{p}=\max_{s,a}\psi_{s,a}$ is the maximal budget. Then $D\leq\frac{2}{(1-\gamma)^{2}}AS^{1+c(q)}\Delta_{p}$ where c(q)=11/qc(q)=1-1/q.
c) Let $\mathcal{P}$ be a non-rectangular $\ell_{q}$ uncertainty set around the nominal $\bar{p}=p_{0}$ such that
$\mathcal{P}=\left\{p:\left\lVert p-\bar{p}\right\rVert_{q}\leq\Delta_{p}\,,\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s,a)=1\quad\forall(s,a)\in\mathcal{S}\times\mathcal{A}\right\}$ (71)
for some $q\geq 1$, where $0<\Delta_{p}$ is the budget. Then $D\leq\frac{2}{(1-\gamma)^{2}}(AS^{2})^{c(q)}\Delta_{p}$, where c(q)=11/qc(q)=1-1/q.
@@
For case a), we have
$\frac{2}{(1-\gamma)^{2}}SA\max_{s\in\mathcal{S},a\in\mathcal{A}}\left\lVert p_{K}(\cdot|s,a)-p_{0}(\cdot|s,a)\right\rVert_{1}$
$\leq\frac{4}{(1-\gamma)^{2}}AS^{2}\,.$
For case b), we have
$\frac{2}{(1-\gamma)^{2}}SA\max_{s\in\mathcal{S},a\in\mathcal{A}}\left\lVert p_{K}(\cdot|s,a)-p_{0}(\cdot|s,a)\right\rVert_{1}$
$\leq\frac{2}{(1-\gamma)^{2}}SAS^{c(q)}\max_{s\in\mathcal{S},a\in\mathcal{A}}\left\lVert p_{K}(\cdot|s,a)-p_{0}(\cdot|s,a)\right\rVert_{q}$
$\leq\frac{2}{(1-\gamma)^{2}}AS^{1+c(q)}\Delta_{p}\,$, (definition of $\Delta_{p}$)
where the second inequality follows via Hölder’s inequality, since for any $X\in\mathbb{R}^{d}$, $\left\lVert X\right\rVert_{1}=\left\lVert\mathbf{1}\cdot X\right\rVert_{1}\leq d^{1/p}\left\lVert X\right\rVert_{q}$ for $p=(1-1/q)^{-1}=c(q)^{-1}$.
For case c), we have
$\frac{1}{(1-\gamma)^{2}}\left|\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\sum_{s^{\prime}\in\mathcal{S}}\left(p_{K}(s^{\prime}|s,a)-p_{0}(s^{\prime}|s,a)\right)\right|+\frac{1}{(1-\gamma)^{2}}\left|\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\sum_{s^{\prime}\in\mathcal{S}}\left(p_{0}(s^{\prime}|s,a)-p_{K}(s^{\prime}|s,a)\right)\right|$
$\leq\frac{2}{(1-\gamma)^{2}}\left\lVert p_{K}-p_{0}\right\rVert_{1}$
$\leq\frac{2}{(1-\gamma)^{2}}(AS^{2})^{c(q)}\left\lVert p_{K}-p_{0}\right\rVert_{q}$
$\leq\frac{2}{(1-\gamma)^{2}}(AS^{2})^{c(q)}\Delta_{p}\,.$ (definition of $\Delta_{p}$)
where the second inequality again makes use of Hölder’s inequality analogous to part b). ∎
@@
Lemma 26 (Lemma 16 of Liu et al. (2021b) rephrased).
@@
For any $\pi\in\Pi,p\in\mathcal{P}$, we have $d_{\rho}^{\pi,p}(s)\geq(1-\gamma)\rho(s)$, such that using derivations as in Theorem 4.2 of Wang et al. (2024),
@@ D.6.2 Main theorem
(a) The proof uses similar derivations as in Theorem 3 of Liu et al. (2021b) and then accounts for the performance difference due to transition kernels.
@@
Filling in the lower bound on the inner product from Eq. 41 from Liu et al. (2021b) (see also Lemma 15),
@@
Applying Lemma 26 to $p^{*}$, it follows that $\left\lVert\lambda^{*}\right\rVert=\mathcal{O}(1)$ such that $\sum_{j=1}^{m}\lambda_{j}^{*}\mathcal{O}(\epsilon)=\mathcal{O}(\epsilon)$. Note then that via Lemma 25, $\frac{\alpha}{1-\gamma}\sum_{k\in[K]}\left(B_{d_{\rho}^{\pi^{*},p_{k-1}}}(\pi^{*},\pi_{k-1})-B_{d_{\rho}^{\pi^{*},p_{k-1}}}(\pi^{*},\pi_{k})\right)\leq\frac{\alpha}{1-\gamma}\left(\log(A)+2\log(\frac{1}{U})\right)$. Moreover, the preconditions for $\alpha$ in the further derivations in Lemma 21 of Liu et al. (2021b) are satisfied. The remainder of the proof removes the constant $\frac{M}{1-\gamma}=\mathcal{O}(1)$ and then follows the derivations in Lemma 21 of Liu et al. (2021b). ∎
@@
D.6.3 Non-rectangular uncertainty sets
The global performance guarantee of CPI can be expressed in terms of the degree of non-rectangularity (Li et al., 2023). Note that we slightly modify the original lemma by removing the irreducibility assumption since the mismatch coefficient can also be limited using Assumption 3.
Lemma 30 (Global performance guarantee of CPI, Theorem 3.8 in Li et al. (2023)).
Let $\epsilon>0$, let $\pi\in\Pi$ and let $V:\mathcal{S}\to\mathbb{R}$ be a value function. Then Algorithm 2 returns within $\mathcal{O}(1/\epsilon^{2})$ iterations a solution $\hat{p}$ that satisfies
$V_{\pi,p^{*}}(\rho)-V_{\pi,\hat{p}}(\rho)\leq M(2\epsilon+\delta_{\mathcal{P}})\,$, (82)
where $\delta_{\mathcal{P}}$ is the degree of non-rectangularity defined in Eq 44.